OpenStudy (anonymous):

Find domain, range, x and y intercept, Horizontal and vertical asymptote? How do I do it? f(x)=3^x-9

OpenStudy (anonymous):

How do you get the x and y int? My solution set says x-in t: (2,0) and y-int: (0,-8). how do you get that answer?

myininaya (myininaya):

to find x intercept let y=0 to find y intercept let x=0

myininaya (myininaya):

to find y intercept; set x=0 so we have y=3^0-9=1-9=-8

OpenStudy (anonymous):

well to get the y-intercept.. just substitute x=0 and solve for y: y=3^0-9 --> y=1-9 -->y=-8, therefor the y-int:(0,-8). to find the x-int, set y=0 and solve for x, that's: 0=3^x-9 --> 3^x=9 --> x=2, so the x-int is (2,0)

myininaya (myininaya):

you beat me :(

OpenStudy (anonymous):

you did half the answer, and I did half.. we are even :)

myininaya (myininaya):

lies

OpenStudy (anonymous):

Thanks guys :)